OpenStudy (anonymous):

x + 5 –1/2 (4 - 6x) = 12 Please Explain and Answer

OpenStudy (anonymous):

x + 5 –(1/2 )(4 - 6x) = 12 like this?

OpenStudy (anonymous):

x+5-1/2-2+3x=12 4x+3and 1/2 =12 -3 and 1/2 on both sides 4x=8.5 divide by four on both sides x=2.125 to check plug in x into the equasion

OpenStudy (anonymous):

I MIGHT HAVE DONE SOMETHING WRONG SO PLZ CHECK